# Quantifying Observational Projection Effects with a Simulation-based hot CGM model

### v1

@@ Abstract
Aims. The hot phase of the circumgalactic medium (CGM) allows us to probe the inflow and outflow of gas within a galaxy, which is responsible for dictating the evolution of the galaxy. Studying the hot CGM sheds light on a better understanding of gas physics, which is crucial to inform and constrain simulation models. With the recent advances in observational measurements probing the hot CGM in X-rays and tSZ, we have a new avenue for widening our knowledge of gas physics and feedback by exploiting the information from current/future observations.
Methods. In this paper, we use the TNG300 hydrodynamical simulations to build a fully self-consistent forward model for the hot CGM. We construct a lightcone and generate mock X-ray observations. We quantify the projection effects, namely the locally correlated large-scale structure in X-rays and the effect due to satellite galaxies misclassified as centrals, which affect the measured hot CGM galactocentric profiles in stacking experiments.
Results. We present an analytical model that describes the intrinsic X-ray surface brightness profile across the stellar and halo mass bins. The increasing stellar mass bins result in decreasing values of $\beta$, the exponent quantifying the slope of the intrinsic galactocentric profiles. We carry forward the current state-of-the-art by also showing the impact of the locally correlated environment on the measured X-ray surface brightness profiles. We also present, for the first time, the effect of misclassified centrals in stacking experiments for three stellar mass bins: $10^{10.5-11}{\rm M}_{\odot}$, $10^{11-11.2}{\rm M}_{\odot}$, and $10^{11.2-11.5}{\rm M}_{\odot}$. We find that the contaminating effect of the misclassified centrals on the stacked profiles increases when the stellar mass decreases. When stacking galaxies of Milky-Way-like stellar mass, this effect is non-negligible already at a low level of contamination: in particular, misclassified centrals contributing 30%, 10%, or 1% of a sample dominate the measured surface brightness profile at radii $\geq 0.06\times R_{200m}$, $\geq 0.2\times R_{200m},$ and $\geq 0.3\times R_{200m}$, respectively.
@@ 1 Introduction
The Circumgalactic medium (CGM) plays a crucial role at interfacing the gas between the interstellar medium (ISM) within galaxies and the external intergalactic medium (IGM) outside them. The CGM is commonly defined as the gas within the virial radii of the galaxy and outside their galactic disks. It is the gravitationally bound gas that encompasses the fossil imprints of the physical mechanisms, such as outflows, inflows, and feedback processes, dictating the evolution of the galaxy (see Tumlinson et al. 2017 for a review). These mechanisms are sensitive to the galaxy’s environment and its halo properties, as shown by the complex Stellar-to-Halo-Mass-Relation (SHMR); see the review from Wechsler & Tinker (2018). According to most models, the SHMR indicates that the low-halo-mass end is sensitive to stellar and supernovae (SN) driven feedback  (Dekel & Woo 2003; Benson et al. 2003), causing heating of the gas, hot bubbles (McKee & Ostriker 1977), galactic wind outflows (Dekel & Silk 1986), and turbulence (Ostriker & Shetty 2011; Strickland & Heckman 2009). In contrast, Active Galactic Nuclei (AGN) should be the main drivers of feedback in the high-halo-mass end (Silk & Rees 1998; Fabian 2012; Eckert et al. 2021). Of particular interest is the peak of the SHMR relation at the pivotal halo mass $M_{\rm h}\sim 10^{12}$ M, similar to our Milky-Way (MW) mass, corresponding to the mass scale where star formation efficiency reaches its maximum. Thus, studying the volume-filling hot gas component of the CGM, especially in the MW halo mass regime, encapsulating a range of physical processes, is crucial for testing galaxy formation models (Faucher-Giguère & Oh 2023).
@@
A plethora of theoretical models (e.g., Faerman et al. 2017; Voit et al. 2019; Pal Choudhury et al. 2019; Stern et al. 2024; Singh et al. 2021; Pandya et al. 2023; Faerman & Werk 2023) have been used to describe and predict the intrinsic properties of the hot CGM; Singh et al. (2024) discuss the comparison of some idealised cases. Similarly, Oppenheimer et al. (2020); Vladutescu-Zopp et al. (2024) made explicit predictions of the intrinsic hot CGM profiles obtained by assigning mock X-ray emission to the hot halos found in cosmological hydrodynamical simulations. However, before testing theoretical and numerical predictions of the hot CGM against state-of-the-art observations, we must quantify all the effects that influence the hot CGM observations in stacking experiments. Among the primary sources of contamination in X-ray stacking experiments, inhibiting us from retrieving the physical properties of the detected CGM emission, are (1) the unresolved AGN and X-ray binaries (XRB) population of galaxies (Biffi et al. 2018; Vladutescu-Zopp et al. 2023), and (2) the projection effects within the hot gas emission of the cosmic web. These projection effects include the contribution from the Large-Scale Structure (LSS) halo environment (locally correlated X-ray emission), the effect of satellite galaxies being misclassified as central due to limitations in the (photometric) redshift accuracy for the galaxies in large surveys (see e.g., Weng et al. 2024), and the Line-of sight (LoS) projection of uncorrelated X-ray emission of fore- and background structures. Z+24 tried to model these effects empirically for the first time; however, we need forward models based on cosmological hydrodynamical simulations for a complete description and full understanding of these contaminating effects.
@@
In this paper, we analyse numerical simulations to empirically quantify the effects of the locally correlated X-ray emission and the misclassified centrals, also called the satellite-boost effect, relevant for measuring the hot CGM in stacking experiments. To do this, we use the TNG300 hydrodynamical simulations (Pillepich et al. 2018a; Marinacci et al. 2018; Naiman et al. 2018; Nelson et al. 2015; Springel et al. 2018) to construct a lightcone and generate mock X-ray observations. We focus here on the projection effects of hot gas affecting the X-ray surface brightness profiles. The structure of the paper is as follows. Sec. 2 details the construction of the TNG300 lightcone (LC-TNG300) used for modelling the hot CGM in this study. We explain all the projection effects in detail in Sec. 2.1. In Sec. 3, we describe the process for generating mock X-ray observations that mimic observational data and provide the machinery for computing the surface brightness profiles. Sec. 4 presents the main results of this work done with the TNG300 lightcone. Lastly, Sec. 5 discusses the main findings and Sec. 6 provides an overview with prospects.
Table 1: For every stellar mass bin, we present the mean stellar mass within twice the stellar half mass radius, the total number of central galaxies, the total number of satellite galaxies, mean halo mass, the minimum and maximum values of halo masses for the given stellar mass bin, the mean $R_{200m}$, and the mean $R_{500c}$.
@@
$10.5-11$ $10.7$ $5109$ $7,826$ $12.7$ $[11.9,13.7]$ $525.02$ $242.75$
$11-11.25$ $11.1$ $680$ $899$ $13.3$ $[12.6,14.0]$ $816.05$ $369.01$
$11.25-11.5$ $11.4$ $305$ $365$ $13.6$ $[13.2,14.2]$ $1077.71$ $484.52$
Table 2: For every halo mass bin used to generate the X-ray surface brightness profiles, we present the mean halo mass, the total number of distinct halos, the mean stellar mass, the mean $R_{\rm 200m}$ and the mean $R_{\rm 500c}$, for every halo mass bin.
@@ 2 Simulated lightcone with IllustrisTNG: LC-TNG300
Figure 2: The projected rest-frame X-ray events from the TNG300 ligthcone in the $0.5-2.0$ keV band for a telescope with energy-independent collecting area $1000$ cm2 and exposure time of $1000$ ks. The events are generated using the hot gas cells within the TNG300 lightcone at redshifts $0.03-0.3$ using pyXsim (ZuHone & Hallman 2016). The centre for the projection onto the sky is chosen as R.A., Dec.(0.,0.$\equiv$(0.,0.)≡ ( 0 . , 0 . ) degrees. The contours represent the two X-ray surface brightness levels of $5\times 10^{-14}$ kev s-1 cm-2 arcmin-2 and $2\times 10^{-16}$ kev s-1 cm-2 arcmin-2.
Figure 3: The projected halos (centrals) are overplotted with their corresponding scaled $R_{500c}$ at $z=0.3$; the $R_{500c}$ of the halos are represented by the size of the circles. We also show the projected rest-frame X-ray events from the TNG300 ligthcone in the $0.5-2.0$ keV band for a telescope with energy-independent collecting area $1000$ cm2 and exposure time of $1000$ ks at the redshift slice of $0.284\leq z\leq 0.3$. The halos with $M_{200m}\in 10^{14-14.5}\ \rm M_{\odot}$ are shown in red, $M_{200m}\in 10^{13.5-14}\ \rm M_{\odot}$ in grey, $M_{200m}\in 10^{13-13.5}\ \rm M_{\odot}$ in cyan, and $M_{200m}\in 10^{12.5-13}\ \rm M_{\odot}$ in blue.
@@
The Friends-of-Friends (FoF) algorithm is applied to the dark matter particles with linking length $b=0.2$ to obtain the halos. The subhalos are retrieved with subfind (Springel et al. 2001; Dolag et al. 2009) that detects gravitationally bound substructures, which are equivalent to galaxies in observations. Additionally, subfind also classified the subhalos into centrals and satellites, where centrals are the most massive substructure within a distinct halo.
@@
2.1 Contributions to observed X-ray surface brightness profiles
@@
    Intrinsic emission from satellites. The X-ray emission from the satellite galaxy contributes to the total CGM emission. As the host galaxy is more massive, with a deeper potential well, the contribution of this effect is negligible in stacking experiments (see Rohr et al. (2024) for further insights on the detectability of the satellite emission).
@@
    Contamination from misclassified centrals. In observations that use photometric surveys, classifications between centrals and satellites are inhibited due to limitations in the redshift accuracy. This effect is large and unavoidable for photometric surveys ($\sim 30\%$ for MW-like galaxies; Sec 3.5 in Z+24) and significantly mitigated for spectroscopic surveys ($\sim 1\%$ for MW-like galaxies; Sec 3.7 in Z+24), but not completely removed due to survey incompleteness or residual uncertainty in the central/satellite classification for systems with a low number of galaxies (see also  Weng et al. 2024 who quantify the effect for absorbers in cold gas with TNG50). This implies that the measured X-ray surface brightness profiles in stacked samples of galaxies classified as central contain the intrinsic emission around truly central galaxies but are contaminated by the emission measured around misclassified centrals. In conclusion, including the emission around satellites in the stacking analysis alters the recovered profiles and, therefore, must be modelled.
@@
This work focuses on modelling the locally correlated environment and the effect due to misclassified-centrals (the satellite-boost effect).
@@ 3 Method
Figure 4: Mean X-ray surface brightness profiles in the stellar mass bins: $M_{\star}=10^{10.5-11}\ {\rm M}_{\odot}$, corresponding to MW-like galaxies (left), $M_{\star}=10^{11-11.2}\ {\rm M}_{\odot}$ (centre), and $M_{\star}=10^{11.2-11.5}\ {\rm M}_{\odot}$ (right). The vertical dashed line is the mean $R_{500c}$ at $242.75$ kpc (left), $369.01$ kpc (centre), and $484.52$ kpc (right) of the respective galaxy stellar mass bin with the shaded area corresponding to the minimum and maximum values. The black line is the analytical model, shown in Eq. 7, fit to the LC-TNG300 mean X-ray surface brightness profiles. The best-fitting parameters for the model are given in Tab. 4. We find decreasing values for the slope, $\beta$, with increasing stellar mass bins.
@@
Table 3: For every stellar mass bins, we present the best-fitting parameters of the model (see Eq. 7): $S_{0}$,:the central surface brightness; $r_{\rm c}$, the core radius; $\beta$, the exponent quantifying the slope of the profile; $r_{s}$, the scale radius at which the slope changes to $\epsilon$.
@@
Table 4: For every halo mass bins, we present the best-fitting parameters of the model (see Eq. 7): $S_{0}$, the central surface brightness; $r_{\rm c}$, the core radius; $\beta$, the exponent quantifying the slope of the profile; $r_{s}$, the scale radius at which the slope changes to $\epsilon$.

### v2

@@ Abstract
Aims. The hot phase of the circumgalactic medium (CGM) allows us to probe the inflow and outflow of gas responsible for dictating the evolution of a galaxy’s structure. Studying the hot CGM sheds light on the physical properties of the gas phase of the baryons, which is crucial to inform and constrain simulation models. With the recent advances in observational measurements probing the hot CGM in X-rays and thermal Sunyaev-Zeldovich (tSZ), we have a new avenue for widening our knowledge of gas physics and feedback.
Methods. In this paper, we use the TNG300 hydrodynamical simulations to build a fully self-consistent forward model for the hot CGM. In order to do that, we construct a lightcone and generate mock X-ray observations of the large-scale structure. We quantify the main projection effects impacting CGM measurements, namely the locally correlated large-scale structure in X-rays and the effect due to satellite galaxies misclassified as centrals, which affect the measured hot CGM galactocentric profiles in stacking experiments.
Results. We present an analytical model that describes the intrinsic X-ray surface brightness profiles of halos across the stellar and halo mass bins. The increasing stellar mass bins result in decreasing values of $\beta$, the exponent quantifying the slope of the intrinsic galactocentric profiles. We measure the effect of misclassified centrals in stacking experiments for three stellar mass bins $10^{10.5-11}{\rm M}_{\odot}$, $10^{11-11.25}{\rm M}_{\odot}$, and $10^{11.25-11.5}{\rm M}_{\odot}$. We find that the contaminating effect of the misclassified centrals on the stacked profiles increases when the stellar mass decreases. When stacking galaxies of Milky-Way-like stellar mass, this effect is dominant already at a low level of contamination: in particular, misclassified centrals contributing 30%, 10%, or 1% of a sample dominate the measured surface brightness profile at radii $\geq 0.11\times R_{500c}$, $\geq 0.24\times R_{500c},$ and $\geq 1.04\times R_{500c}$, respectively.
@@ 1 Introduction
The circumgalactic medium (CGM) plays a crucial role in interfacing the gas between the interstellar medium (ISM) within galaxies and the external intergalactic medium (IGM) outside them. The CGM is commonly defined as the gas within the virial radii of the galaxy and outside their stellar disks. It is the gravitationally bound gas that encompasses the fossil imprints of the physical mechanisms, such as outflows, inflows, and feedback processes, dictating the evolution of the galaxy (see Tumlinson et al. 2017 for a review). These mechanisms are sensitive to the galaxy’s environment and its halo properties, as shown by the complex Stellar-to-Halo-Mass-Relation (SHMR); see the review from Wechsler & Tinker (2018). According to most models, the SHMR indicates that the low-halo-mass end is sensitive to stellar and supernovae (SN) driven feedback  (Dekel & Woo 2003; Benson et al. 2003), causing heating of the gas, hot bubbles (McKee & Ostriker 1977), galactic wind outflows (Dekel & Silk 1986), and turbulence (Ostriker & Shetty 2011; Strickland & Heckman 2009). In contrast, Active Galactic Nuclei (AGN) should be the main drivers of feedback in the high-halo-mass end (Silk & Rees 1998; Fabian 2012; Eckert et al. 2021). Of particular interest is the peak of the SHMR relation at the pivotal halo mass $M_{\rm h}\sim 10^{12}$ M, similar to our Milky-Way (MW) mass, corresponding to the mass scale where star formation efficiency reaches its maximum. Thus, studying the volume-filling hot gas component of the CGM, especially in the MW halo mass regime, encapsulating a range of physical processes, is crucial for testing galaxy formation models (Faucher-Giguère & Oh 2023).
@@
A plethora of theoretical models (e.g., Faerman et al. 2017; Voit et al. 2019; Pal Choudhury et al. 2019; Stern et al. 2024; Singh et al. 2021; Pandya et al. 2023; Faerman & Werk 2023) have been used to describe and predict the intrinsic properties of the hot CGM; Singh et al. (2024) discuss the comparison of some idealised cases. Similarly, Oppenheimer et al. (2020); Vladutescu-Zopp et al. (2025) made explicit predictions of the intrinsic hot CGM profiles obtained by assigning mock X-ray emission to the hot halos found in cosmological hydrodynamical simulations. However, before testing theoretical and numerical predictions of the hot CGM against state-of-the-art observations, we must quantify all the effects that influence the hot CGM observations in stacking experiments. Among the primary sources of contamination in X-ray stacking experiments, inhibiting us from retrieving the physical properties of the detected CGM emission, are (1) the unresolved AGN and X-ray binaries (XRB) population of galaxies (Biffi et al. 2018; Vladutescu-Zopp et al. 2023), and (2) the projection effects within the hot gas emission of the cosmic web. These projection effects include the contribution from the Large-Scale Structure (LSS) halo environment (locally correlated X-ray emission), the effect of satellite galaxies being misclassified as central due to limitations in the (photometric) redshift accuracy for the galaxies in large surveys (see, e.g., Weng et al. 2024), the offset between the X-ray centre and the centre defined by the minimum of the halo potential, and the Line-of sight (LoS) projection of uncorrelated X-ray emission of fore- and background structures. Z+24 tried to model these effects empirically for the first time; however, we need forward models based on cosmological hydrodynamical simulations for a complete description and full understanding of these contaminating effects.
@@
In this paper, we analyse numerical simulations to empirically quantify the effects of the locally correlated X-ray emission and the misclassified centrals, also called the satellite-boost effect, relevant for measuring the hot CGM in stacking experiments. To do this, we use the TNG300 hydrodynamical simulations (Pillepich et al. 2018a; Marinacci et al. 2018; Naiman et al. 2018; Nelson et al. 2015; Springel et al. 2018) to construct a lightcone and generate mock X-ray observations. We focus here on the projection effects of hot gas affecting the X-ray surface brightness profiles. The structure of the paper is as follows. Sec. 2 details the construction of the TNG300 lightcone (LC-TNG300) used for modelling the hot CGM in this study. We explain all the projection effects in detail in Sec. 3. In Sec. 4, we describe the process for generating mock X-ray observations that mimic observational data and provide the machinery for computing the surface brightness profiles. Sec. 5 presents the main results of this work done with the TNG300 lightcone. Lastly, Sec. 6 discusses the main findings and Sec. 7 provides an overview with prospects.
Table 1: Stellar mass bins used to generate the X-ray surface brightness profiles.
@@
$10.5-11$ $10.7$ $5109$ $2719$ $12.7$ $[11.9,13.7]$ $525.02$ $242.75$
$11-11.25$ $11.1$ $680$ $220$ $13.3$ $[12.6,14.0]$ $816.05$ $369.01$
$11.25-11.5$ $11.4$ $305$ $60$ $13.6$ $[13.2,14.2]$ $1077.71$ $484.52$
 [^1] We present the mean stellar mass within twice the stellar half mass radius, the total number of central galaxies, the total number of satellite galaxies, mean halo mass, the minimum and maximum values of halo masses for the given stellar mass bin, the mean $R_{200m}$, and the mean $R_{500c}$.
Table 2: Halo mass bins used to generate the X-ray surface brightness profiles.
@@
 [^2] We present the mean halo mass, the total number of distinct halos, the mean stellar mass, the mean $R_{\rm 200m}$ and the mean $R_{\rm 500c}$.
@@ 2 Simulated lightcone with IllustrisTNG: LC-TNG300
Figure 2: Projected rest-frame X-ray events from the TNG300 ligthcone in the $0.5-2.0$ keV band for a telescope with energy-independent collecting area $1000$ cm2 and exposure time of $1000$ ks. The events are generated using the hot gas cells within the TNG300 lightcone at redshifts $0.03-0.3$ using pyXsim (ZuHone & Hallman 2016). The centre for the projection onto the sky is chosen as R.A., Dec.(0.,0.$\equiv$(0.,0.)≡ ( 0 . , 0 . ) degrees. The contours represent the two X-ray surface brightness levels of $5\times 10^{-14}$ erg s-1 cm-2 arcmin-2 and $2\times 10^{-16}$ erg s-1 cm-2 arcmin-2.
Figure 3: Projected halos (centrals) are overplotted with their corresponding scaled $R_{500c}$ at $z=0.3$; the $R_{500c}$ of the halos are represented by the size of the circles. We also show the projected rest-frame X-ray events from the TNG300 ligthcone in the $0.5-2.0$ keV band for a telescope with energy-independent collecting area $1000$ cm2 and exposure time of $1000$ ks at the redshift slice of $0.284\leq z\leq 0.3$. The halos with $M_{200m}\in 10^{14-14.5}\ \rm M_{\odot}$ are shown in red, $M_{200m}\in 10^{13.5-14}\ \rm M_{\odot}$ in grey, $M_{200m}\in 10^{13-13.5}\ \rm M_{\odot}$ in cyan, and $M_{200m}\in 10^{12.5-13}\ \rm M_{\odot}$ in blue.
@@
The Friends-of-Friends (FoF) algorithm is applied to the dark matter particles with linking length $b=0.2$ to obtain the halos. The subhalos are retrieved with subfind (Springel et al. 2001; Dolag et al. 2009), which detects gravitationally bound substructures that are equivalent to galaxies in observations. Additionally, subfind also classified the subhalos into centrals and satellites, where centrals are the most massive substructure within a distinct halo.
@@
[…]
3 Contributions to observed X-ray surface brightness profiles
@@
    Intrinsic emission from satellites. The X-ray emission from the satellite galaxy contributes to the total observed CGM emission in projection. As the host galaxy is more massive, with a deeper potential well, the contribution of this effect is negligible in stacking experiments (see Rohr et al. (2024) for further insights on the detectability of the satellite emission).
@@
    Contamination from misclassified centrals. In observations that use photometric surveys, classifications between centrals and satellites are inhibited due to limitations in the redshift accuracy. This effect is large and unavoidable for photometric surveys ($\sim 30\%$ contaminants for MW-like galaxies; Sec 3.5 in Z+24) and significantly mitigated for spectroscopic surveys ($\sim 1\%$ contaminants for MW-like galaxies; Sec 3.7 in Z+24), but not completely removed due to survey incompleteness or residual uncertainty in the central/satellite classification for systems with a low number of galaxies (see also  Weng et al. 2024 who quantify the effect for absorbers in cold gas with TNG50). This implies that the measured X-ray surface brightness profiles in stacked samples of galaxies classified as centrals contain the intrinsic emission around truly central galaxies but are contaminated by the emission measured around satellites (off-centred) misclassified as centrals. In conclusion, including the emission around satellites in the stacking analysis alters the recovered profiles and, therefore, must be modelled.
@@
    The X-ray to true centre offset. The theoretically defined centre of a halo, which is the minimum of the dark matter potential well, could be offset from the observationally defined centres, e.g., the peak of the SZ signal or the X-ray emission, or the optical/infrared centre inferred from the stellar distribution  (Zitrin et al. 2012; Lauer et al. 2014; Saro et al. 2015; Cui et al. 2016; Seppi et al. 2023). This offset is associated with physical effects such as galaxy mergers (Martel et al. 2014), misidentification of the brightest halo galaxy (Hoshino et al. 2015; Hikage et al. 2018; Oguri et al. 2018; Zhang et al. 2019) or failure of the brightest halo galaxy being a proxy for the minimum of the halo potential in optical studies (Skibba et al. 2011). Overall, miscentering causes a smoothing effect in stacking studies (Oguri & Takada 2011), and therefore, must be calibrated and modelled. This effect of miscentering correlates with halo mass; hence, it is well-studied at the cluster and group scales. As we simulate X-ray emission around lower mass, MW-like halos in this work, we extend the discussion of miscentering of the X-ray emission from the minimum of the host halo potential to MW mass scales (see discussion in Sec. 6.3).
  6. 6.
@@
  8. 8.
    Other X-ray background components. The local flat background is a combination of (1) the Soft X-ray background (SXRB; Henley & Shelton 2010, 2012, 2013; Nakashima et al. 2018; Gupta et al. 2023; Pan et al. 2024; Yeung et al. 2024), which is composed of emission from the local hot bubble, the MW CGM, other intervening galactic structure and the time variable solar-wind charge exchange, (2) the Cosmic X-ray background (De Luca & Molendi 2004; Gilli et al. 2007; Luo et al. 2016; Cappelluti et al. 2017), which is dominated by AGN, and (3) the instrumental effects. In this study, we simulate intrinsic X-ray events from hot gas alone; therefore, this study does not expand on the impacts of the other well-studied X-ray background components.
This work focuses on modelling the locally correlated environment and the effect due to misclassified-centrals (the satellite-boost effect). We also quantify the effect of the X-ray to true centre offset for the halo mass bins considered in this work.
@@ 4 Method
Figure 4: Mean X-ray surface brightness profiles in the stellar mass bins: $M_{\star}=10^{10.5-11}\ {\rm M}_{\odot}$, corresponding to MW-like galaxies (left), $M_{\star}=10^{11-11.25}\ {\rm M}_{\odot}$ (centre), and $M_{\star}=10^{11.25-11.5}\ {\rm M}_{\odot}$ (right). The vertical dashed line is the mean $R_{500c}$ at $242.75$ kpc (left), $369.01$ kpc (centre), and $484.52$ kpc (right) of the respective galaxy stellar mass bin with the shaded area corresponding to the minimum and maximum values. The black line is the analytical model, shown in Eq. 7, fit to the LC-TNG300 mean X-ray surface brightness profiles. The best-fitting parameters for the model are given in Tab. 4. We find decreasing values for the slope, $\beta$, with increasing stellar mass bins.
@@
Table 3: For every stellar mass bin, we present the best-fitting parameters of the model (see Eq. 7).
@@
Table 4: For every halo mass bin, we present the best-fitting parameters of the model (see Eq. 7).
@@
777: $S_{0}$, the central surface brightness; $r_{\rm c}$, the core radius; $\beta$, the exponent quantifying the slope of the profile; $r_{s}$, the scale radius at which the slope changes to $\epsilon$.